A random sample of size 100 results in a sample mean of 100. The population standard deviation is known to be 35. What is the value of the standard error of the mean?


1
Expert's answer
2021-08-11T13:23:08-0400

"n=100 \\

\mu=100 \\

\sigma=35 \\

SE= \frac{\sigma}{\sqrt{n}} \\

= \frac{35}{\sqrt{100}} \\

= 3.5"